OpenStudy (amazingmeisnotonfire):

Will give medal 15=8x^2-14x Solve by factoring

OpenStudy (welshfella):

8x^2 - 14x - 15 = 0 factors of 8 would be 4 * 2 OR 8 * 1 factors of -15 are -3 * 5 or 3 * -5 or -15 * 1 or 15 * -1 we have to combine one of the first set to one of the second to get a result of -14 when they are added

OpenStudy (welshfella):

this one is tricky

OpenStudy (amazingmeisnotonfire):

Still confuzzled, but I found a different way to do it and the answers are multiple choice and it's one of them, so I'm gonna go with it.

OpenStudy (welshfella):

its -5 and 4 with the 3 and the 2

OpenStudy (welshfella):

-20 + 6 = -14

OpenStudy (welshfella):

(2x - 5 )(4x + 3) = 0

OpenStudy (amazingmeisnotonfire):

The answers I got were -3/4 and 5/2

OpenStudy (welshfella):

yes thats correct

OpenStudy (amazingmeisnotonfire):

K then

OpenStudy (welshfella):

when you have coefficients that have a few different factors these are tricky

OpenStudy (welshfella):

which wat did you do it ?

OpenStudy (welshfella):

there is one method called the ac method where you multiply the first coefficient by theb last and then check the factors here ac = 8 * -15 = -120

OpenStudy (welshfella):

i prefer just to play around with the factors

OpenStudy (anthonyym):

15 = 8x^2 - 14x 0 = 8x^2 - 14x - 15 Now you need two numbers that add to -14 and multiply to the product of 8 and -15, which is -120. I like to organize it like this The numbers are -20 and 6. Now break down -14x into -20x + 6x 0 = 8x^2 - 20x + 6x - 15 0 = 4x(2x - 5) + 3(2x -5) Factor out 2x-5 0 = (2x - 5)(4x + 3) 2x - 5=0 x = 5/2 4x + 3 = 0 x = -3/4

OpenStudy (welshfella):

yes thats the ac method